OpenStudy (anonymous):

A football field appears to be flat, its surface is actually shaped like a parabola so that rain runs off to both sides. The cross section of a field with synthetic turf can be modeled by y = -0.000234x(x - 160) where x and y are measured in feet. a) What is the fields length? b) What is the maximum height of the field's surface? @whpalmer4

OpenStudy (whpalmer4):

\[y = -0.000234x(x-160)\]I guess the x axis runs the length of the field, so one end is at x = 0, and the other end is wherever y = 0 (other than at x =0). Do you agree?

OpenStudy (anonymous):

@whpalmer4 yes

OpenStudy (anonymous):

@whpalmer4 well ok so this question comes with a graph do you want me to upload it?

OpenStudy (anonymous):

@whpalmer4 i don't know if it helps you

OpenStudy (whpalmer4):

Yes, that's what I was envisioning. Okay, so how do we find the field's length?

OpenStudy (whpalmer4):

The length of the field is the distance between the two points where the curve crosses the x-axis, right?

OpenStudy (whpalmer4):

If we plug in \(x = 0\), we have \[y = -0.000234(0)(0-160) = 0\]

OpenStudy (anonymous):

@whpalmer4 ok so how would i find the maximum height of the field's surface?

OpenStudy (whpalmer4):

Let's expand the formula: \[y = -0.000234x^2 + 0.03744x\]\[0 = -0.000234x^2+0.03744x\]We already found \(x=0\) as one solution, but we need the other. Probably the easiest thing here is to use the quadratic formula: \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]which gives the solutions to \[ax^2+bx+c=0,\,a\ne0\] We have \[a = -0.000234,b = 0.03744,c=0\] \[x = \frac{-0.03744\pm\sqrt{0.03744^2-4(-0.000234)(0)}}{2(-0.000234)}\]which simplifies to \[x=\frac{-0.03744\pm 0.03744}{-0.000468}\]

OpenStudy (whpalmer4):

I think you can get the non-zero value of \(x\) from that, right? Now to find the maximum height, which is at the vertex of the parabola. There are a number of ways we could do this. One is to remember that a parabola is symmetric about its vertex, and so the vertex will be centered between the two values of \(x\) that make \(y=0\). Divide the length of the field by 2, in other words, and that's the x-coordinate of the vertex. Plug that value of \(x\) into the formula for the height of the cross-section, and you've got your maximum height.

OpenStudy (whpalmer4):

Now, if you don't remember that the vertex is centered between the two x-intercepts, you might remember that the x-coordinate of the vertex where the equation is written in the form \(y = ax^2 + bx + c\) is at \[x = -\frac{b}{2a}\]which would give us an x-value of \[x=-\frac{0.03744}{-0.000468} = \] yet another way would be to use the "vertex form" of a parabola, which is \[y = a(x-h)^2+k\]with the vertex at \((h,k)\)Putting our equation in comparison: \[y=a(x-h)^2+k\]\[ y = -0.000234x^2 + 0.03744x\]we would have to complete the square to make that work. Let's not, because we'd just get the same answer for a bunch of extra work :-)

OpenStudy (whpalmer4):

Actually, looking at the problem, we could have done this more easily. \[y = -0.000234x(x-160)\]The solutions aren't affected at all by the -0.000234 out front: that only affects how high the parabola goes (or how low), but not where it crosses the x-axis! Really, we just need to find the solutions to \[0 = x(x-160)\]And that's easy!\[0=x(x-160)\]is solved by \[x=0\]\[x-160 = 0\]so our solutions are \[x = 0, 160\] Half the length of the field is the midpoint, at x = 80. The height there is \[-0.000234(80)(80-160) = -0.000234*-6400 = 1.4976\]

OpenStudy (whpalmer4):

My eyes just glazed over when I saw the big ugly decimal out front, and kept me from spotting the easier path to the solution :-)

OpenStudy (anonymous):

@whpalmer4 I'm sorry you had to go through all of this

OpenStudy (anonymous):

@whpalmer4 Thank you you're the best :D

OpenStudy (whpalmer4):

Considering what you had for the prior two problems, I should have tried to see it in the same fashion. No problem, a little extra work is good for me :-)

OpenStudy (whpalmer4):

Did I succeed in thoroughly confusing you?

OpenStudy (anonymous):

@whpalmer4 Yes you did i was lost :O i was going to ask you if there was another way you could of help me because i was in confusion haha thank you so much :D

OpenStudy (whpalmer4):

Let me try to make it as clear as possible, now that I've made it as unclear as possible :-) \[y = -0.000234x(x-160)\] is the height of the field at any point along its length. The ends of the field are the two points where the height is 0. \[y=-0.000234x(x-160)\]is 0 when either \[-0.000234x = 0\]which happens at \(x =0\)or when \[(x-160) = 0\]which happens when \[x = 160\] So, the length of the field is the distance between x = 0 and x = 160, which is 160. Clear?

OpenStudy (anonymous):

@whpalmer4 ok so a) is 160 and b) is 1.4976 right?

OpenStudy (whpalmer4):

Now, the formula is that of a parabola, which has its vertex equidistant between the two x-intercepts. That means the vertex has an x-coordinate of \[x=\frac{0+160}{2} = 80\]Now we just find the height of the field at \(x=80\) and we are all done: \[y = -0.000234(80)(80-160) = 1.4976\] What we are actually looking at is the distance across the football field (from side to side), not the length of the football field. Problem is written a bit misleadingly, in my opinion. At first I was wondering what kind of an idiot would think a football field is only 160 feet long :-)

OpenStudy (whpalmer4):

Yes, a is 160, b is 1.4976. Probably 1.5 is close enough, considering it's built on a pile of dirt :-)

OpenStudy (whpalmer4):

I hope that cleared it up, I've got to go!

OpenStudy (anonymous):

@whpalmer4 Thank you so much you didn't have to explain all of this thank you for taking the time to bare with me i appreciate it :D <3 :*